OpenStudy (anonymous):

Please help! Will fan and medal! Don't want answer, just want some to explain how to get the answer! Thank you! :) Solve: 6C3 a. 120 b. 720 c. 20 d. 240

OpenStudy (solomonzelman):

C(n,r) = C(6,3) = 6! / (3!(6 - 3)!) = 6! / (3!(3)!) = (1×2×3×4×5×6) / (1×2×3) ( (1×2×3) = (4×5×6) / ( (1×2×3) = (cancel out some terms on top and bottom... 2×5×2 = 20

OpenStudy (anonymous):

okay thanks! :)